Moody Corporation uses a job-order costing system with a plantwide predetermined overhead rate based on machine-hours. At the beginning of the year, the company made the following estimates: Machine-hours required to support estimated production 100,000 Fixed manufacturing overhead cost Variable manufacturing overhead cost per machine-hour $3.ee $650,0e0 Required 1. Compute the plantwide predetermined overhead rate. 2. During the year, Job 400 was started and completed. The following information was available with respect to this job: Direct materials Direct labor cost Machine-hours used $450 $210 40 Compute the total manufacturing cost assigned to Job 400 3. If Job 400 includes 52 units, what is the unit product cost for this job? 4. If Moody uses a markup percentage of 120% of its total manufacturing cost, then what selling price per unit would it have established for Job 400?
Answer #1

Answer

  • All working forms part of the answer
  • Requirement 1

A

Fixed manufacturing overhead cost

$              650,000

B

Machine hours required

                 100,000

machine hours

C = A/B

Fixed overhead rate

$                     6.50

per machine hour

D

Variable overhead rate

$                     3.00

per machine hour

E = C+D

Plantwide Predetermined Overhead rate

$                     9.50 = Answer

per machine hour

  • Requirement 2

A

Machine hours used

40

B

Plantwide Predetermined Overhead rate

$                     9.50

C = A x B

Total manufacturing cost assigned to Job 400

$                      380 = Answer

  • Requirwement 3

A

Direct materials

$                      450

B

Direct labor cost

$                      210

C

Total manufacturing cost assigned to Job 400

$                      380

D = A+B+C

Total cost for the Job

$                  1,040

E

Total units for Job 400

52

F = D/E

Units product cost for Job 400

$                        20 per unit = Answer

  • Requirement 4

A

Units product cost for Job 400

$                        20

B = A x 120%

Mark up at 120%

$                        24

C = A+B

Selling price per unit

$                        44 per unit = Answer